The flywheel of a steam engine runs with a constant angular velocity of 180 rev/min. When steam is shut off, the friction of the bearings and of the air stops the wheel in 2.6 h. (a) What is the constant angular acceleration, in revolutions per minute-squared, of the wheel during the slowdown? (b) How many revolutions does the wheel make before stopping? (c) At the instant the flywheel is turning at 90.0 rev/min, what is the tangential component of the linear acceleration of a flywheel particle that is 31 cm from the axis of rotation? (d) What is the magnitude of the net linear acceleration of the particle in (c)?
Answer #1

here,

angular speed , w0 = 180 rev/min = 18.84 rad/s

time , t = 2.6 h = 9360 s

a)

let the constant accelration be alpha

w = w0 + alpha * t

0 = 18.84 + alpha * 9360

alpha = 2.01 * 10^-3 rad/s^2

alpha = 1.15 rev/min^2

b)

angle travelled , theta = w0^2 /( 2 * alpha)

theta = (18.84)^2 /( 2 * 0.00201)

theta = 88294 rad

the number of revolutions , n = theta /2pi = 14059 rev

c)

the tangential linear accelration , at = 0.31 * 0.00201 = 6.51 * 10^-4 m/s^2

d)

w = 90 rev/min = 9.42 rad/s

ac = w^2 * r

ac = 9.42^2 * 0.31 = 27.5 m/s^2

the magnitude of net linear accelration , a = sqrt(ac^2 + at^2)

a = 27.5 m/s^2